### v1

# Effort: Efficient Orthogonal Modeling for Generalizable AI-Generated Image Detection

Abstract
Existing AI-generated image (AIGI) detection methods often suffer from limited generalization performance. In this paper, we identify a crucial yet previously overlooked asymmetry phenomenon in AIGI detection: during training, models tend to quickly overfit to specific fake patterns in the training set, while other information is not adequately captured, leading to poor generalization when faced with new fake methods. A key insight is to incorporate the rich semantic knowledge embedded within large-scale vision foundation models (VFMs) to expand the previous discriminative space (based on forgery patterns only), such that the discrimination is decided by both forgery and semantic cues, thereby reducing the overfitting to specific forgery patterns. A straightforward solution is to fully fine-tune VFMs, but it risks distorting the well-learned semantic knowledge, pushing the model back toward overfitting. To this end, we design a novel approach called Effort: Efficient orthogonal modeling for generalizable AIGI detection. Specifically, we employ Singular Value Decomposition (SVD) to construct the orthogonal semantic and forgery subspaces. By freezing the principal components and adapting the residual components ($\sim$0.19M parameters), we preserve the original semantic subspace and use its orthogonal subspace for learning forgeries. Extensive experiments on AIGI detection benchmarks demonstrate the superior effectiveness of our approach.
Equal Contribution. Corresponding Author.
Figure 1: Evidence for illustrating the asymmetry phenomenon in AI-generated image detection. We show that the baseline detector (i.e., Xception [70]) tends to quickly overfit to the fake patterns in the training set, causing the generalization issue when facing previously unseen fake methods.
1 Introduction
The rapid development of AI generative technologies has significantly lowered the barrier for creating highly realistic fake images. As deep generative models advance and mature [21, 28, 68], the proliferation of AI-generated images (AIGIs) has drawn considerable attention, driven by their ability to produce high-quality content with relative ease. However, these advancements also introduce significant risks, if misused for malicious purposes such as deepfakes (mainly including face-swapping [39] and face-reenactment [81]), which may violate personal privacy, spread misinformation, and erode trust in digital media. Consequently, there is an urgent need to develop a reliable and robust framework for detecting AIGIs. [^1] In the context of this research, AIGI primarily refers to deepfakes (face-swapping) and synthetic images (e.g., nature or arts).
Most existing studies in AIGI detection [86, 70] typically approach the real/fake classification problem as a symmetric binary classification task, akin to the “cat versus dog” problem. A standard binary classifier, often based on deep neural networks, is trained to distinguish between real and fake images by predicting the likelihood of a given test image being fake during inference. Although this paradigm yields promising results when the training and testing distributions (in terms of fake generation methods) are similar, its performance tends to degrade significantly when applied to previously unseen fake methods.
Figure 2: Illustration of the comparison without and with incorporating semantic knowledge. Naively trained models rely solely on forgery patterns for discrimination (the left), while incorporating semantic knowledge expands the discrimination space (the right), such that the discrimination is decided by both forgery and semantic cues, thereby alleviating the overfitting to seen fakes.
To understand the underlying reasons for the failure in generalization, we conducted extensive preliminary investigations and identified an asymmetry phenomenon in AIGI detection: naively trained models tend to take the shortcut and quickly overfit the specific fake patterns in the training set, resulting in the undesirable generalization performance when facing novel fake methods. Visualization in Fig. 1 corroborates this observation. Specifically, the vanilla detector (i.e., Xception [70]) quickly fits the fake patterns at the very early training stage (only a few iterations), resulting in a very low loss of fake, while the real loss is significantly higher than the fake loss ($\sim$ 100$\times$ larger). This is likely because existing AIGI detection datasets [70, 86] typically contain limited and homogeneous fake types, while real samples exhibit significantly greater diversity and variance (such as different categories and scenarios). This asymmetry causes the detection model to quickly overfit to the seen forgery patterns during training, limiting its generalization to detect unseen fakes.
To address this, a key insight is to incorporate the rich semantic knowledge embedded within the large-scale vision foundation models (VFMs) to expand the previous discriminative space (relying on forgery pattern only), such that the discrimination is decided by both forgery and semantic cues. We provide an illustration example in Fig. 2 for intuitively understanding the role of semantic knowledge. Specifically, naively trained models rely solely on forgery patterns for discrimination (the left), fitting well to the seen fakes but failing to generalize previously unseen fakes. When semantic knowledge is incorporated, the discrimination space is largely expanded, with different semantic categories such as Bird and Fish separated (the right). In this case, the discrimination is determined by both forgery and semantic cues, thus reducing the overfitting to the seen forgery patterns and enhancing the model’s generalization.
So, is naively fine-tuning a VFM sufficient to obtain a generalizable AIGI detector? Unfortunately, it risks distorting the well-learned semantic knowledge of the original VFM, pushing the VFM model back toward overfitting. The t-SNE results in Fig. 3 show that fully fine-tuning a VFM can still overfit the seen fake patterns, due to the mentioned asymmetry problem. Specifically, both the Xception (Vanilla CNN) and CLIP [65] (VFM) detector group only the specific fake patterns within the training set into a single cluster, while all other data, including real and fake from unseen manipulation methods are classified into the other cluster. This highlights that simply fine-tuning the VFMs (even using CLIP) is still insufficient to overcome the asymmetry issue and achieve robust generalization.
Figure 3: t-SNE visualizations between Xception and CLIP (full-finetune). We show that both models only learn the specific fake patterns within the training set, treating samples with seen fake patterns as fake while other samples are all considered real, thereby limiting their generalization in detecting unseen fakes.
To this end, we design a novel approach called Effort: Efficient orthogonal modeling for generalizable AIGI detection. Specifically, we employ Singular Value Decomposition (SVD) to construct the orthogonal semantic and forgery subspaces. By freezing the principal components and adapting the residual components ($\sim$0.19M parameters), we preserve the original semantic subspace and use its orthogonal subspace for learning forgeries. We conduct experiments on both deepfake detection and synthetic image detection benchmarks and surprisingly find that our approach achieves significant superiority over other SOTAs with very few trainable parameters.
Overall, this work makes the following key contributions:
  • Asymmetry Phenomenon in AIGI Detection: We introduce the concept of asymmetry in AIGI detection, which has often been overlooked in previous works. A naively trained detector tends to quickly fit the seen fake methods well but, in doing so, it often overfits to specific fake patterns in the training set, limiting its generalization ability to detect unseen fake methods.
  • Semantic Knowledge Involved and Orthogonality: We propose a novel approach, Effort, to address the asymmetry in AIGI detection by (1) leveraging the rich semantic knowledge within the Vision Foundation Models (VFMs) to explain the previous discrimination space, alleviating the over-relying on forgery patterns only, and (2) Using SVD to construct the orthogonal semantic and forgery subspaces, avoiding the distortion of original well-learned semantic knowledge during learning fakes.
  • Stricking Result and Efficiency: Effort achieves remarkably high generalization performance on previously unseen forgeries using nearly only 0.19M tunable parameters, making it efficient and potentially scalable. Moreover, our approach can be applied to detect a broad range of AIGIs, including but not limited to face forgeries (e.g., classical deepfakes) and synthetic natural images.
2 Related Work
Our work focuses on detecting AI-generated images (AIGIs), especially deepfake images (e.g., face-swapping) and synthetic images (e.g., nature or art), following [94]. As the majority of recent works specifically focus on dealing with the generalization issue, where the training and testing distribution differ (in terms of fake methods), we will briefly introduce the classical and recent detection methods toward generalization in deepfake and synthetic images, respectively.
Generalizable Deepfake Image Detection.
The task of deepfake detection grapples profoundly with the issue of generalization. To tackle the generalization issue, one mainstream approach is fake pattern learning. Most existing works are within this line. These methods generally design a “transformation function”, e.g., frequency transformation [42, 53, 47], blending operations [44, 99, 72, 8], reconstruction [104, 4], content/ID disentanglement [91, 30, 20], to transform the original input $x$ into $x^{\prime}$, where they believe that the more general fake patterns can be captured within the feature space of $x^{\prime}$ compared to $x$. However, given the ever-increasing diversity of forgery methods in the real world, it is unrealistic to elaborate all possible fake patterns and “expect” good generalization on unseen fake methods. Another notable direction is to real distribution learning, with a specific methodology involved: one-class anomaly detection [35, 40]. Specifically, [35] introduced a one-class-based anomaly detection, where “abnormal” data is detected by the proposed reconstruction error as the anomaly score. [40] proposed a similar approach to create pseudo-fake “anomaly” samples by using image-level blending on different facial regions. However, it is challenging to ensure that the detector can learn a robust representation of real images by using the very limited real data in existing deepfake datasets (e.g., the FF++ dataset [70] contains only 1,000 real videos with imbalanced facial attribute distributions [83]).
Generalizable Synthetic Image Detection.
With the rapid advancement of existing AI generative technologies, the scope of forged content has expanded beyond facial forgeries to encompass a wide range of scenes. In this context, similar to the deepfake detection field, most existing works typically focus on fake pattern learning that mines the low-level forgery clues from different aspects. Specifically, several approaches have been proposed to capture low-level artifacts, including RGB data augmentations [86], frequency-based features [31], gradients [76], reconstruction artifacts [87, 6, 54], and neighboring pixel relationships [79], random-mapping feature [77]. To illustrate, BiHPF [31] amplifies artifact magnitudes through the application of dual high-pass filters, while LGrad [76] uses gradient information from pre-trained models as artifact representations. NPR [79] introduces a straightforward yet effective artifact representation by rethinking up-sampling operations. In addition to learning from scratch, there are also several research works [61, 89, 48] that perform fake pattern learning by leveraging the existing vision foundation models. For instance, UniFD [61] directly freezes the visual encoder of the pre-trained CLIP model and tunes only a linear layer for binary classification, demonstrating effective deepfake detection even with previously unseen sources. LASTED [89] proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, advancing the field of synthetic image detection. These arts have shown notable improvement in generalization performance when facing previously unseen fake methods.
Figure 4: The proposed SVD-based adaptation approach for AIGI detection. The left branch is the decomposition matrix of the principle components approximation using SVD, while the right residual branch enables the orthogonal learning of real/fake discriminative features.
3 Methodology
3.1 Motivation
Detecting AI-generated images (AIGI) has become an increasingly important topic. Existing works typically formulate AIGI detection as a standard binary classification problem (just like “cat versus dog”) and often get unsatisfied generalization performance. Our preliminary investigation surprisingly found that the failure of generalization can be attributed to the previously overlooked asymmetry phenomenon, where a naive trained detector tends to quickly overfit the specific fake pattern present in the training set, thereby limiting its generalization when facing previously unseen fake methods. To better understand the reasons behind the failure generalization, we present three important research questions (RQs) below.
RQ-1: Why do naively trained detectors fail to generalize to new unseen fake methods?
In the previous sections, we have shown that the asymmetry phenomenon, where the naively trained models tend to quickly overfit the specific fake patterns in the training set, resulting in undesirable generalization performance when facing unseen fake methods (see Fig. 1). This is likely because existing AIGI detection datasets [70, 86] typically contain limited and homogeneous fake types, while real samples exhibit significantly greater diversity and variance. This asymmetry causes the detection model to quickly overfit to the specific and limited forgery methods during training, dominating the model’s discrimination space and shaping the whole space to be low-ranked (see Fig. 3(a)), thereby limiting its representational capacity and eventually hurting the generalization. We will provide a deeper analysis and evidence of this claim in Sec. 4.3.
RQ-2: Why incorporating semantic knowledge helps to alleviate the model’s overfitting?
Given RQ-1, to address the asymmetric problem, a key insight is to leverage the rich semantic knowledge of VFMs to expand and enrich the low-ranked discrimination space (relying on forgery patterns only) to higher-ranked, encouraging the final discrimination based on both forgery patterns and semantic knowledge, thereby improving the alleviating the overfitting to the forgery patterns. Recall that there are numerous well-trained and generalizable vision foundation models (VFMs) available such as CLIP [65], SigLIP [96], and BEIT-v2 [63], which have already learned rich and robust semantic representation. Thus, leveraging the extensive semantic knowledge within VFMs presents a reasonable solution to expand the low-ranked discrimination space and reduce overfitting.
RQ-3: Why do naively fine-tuning VFMs fail to generalize to new unseen fake methods?
A straightforward solution that incorporates semantic knowledge into detection is fine-tuning VFMs, but it risks distorting the well-learned semantic knowledge, falling back to the overfitting to the forgery patterns, similar to a vanilla CNN baseline, as shown in Fig. 3. To overcome these issues, we propose to achieve the orthogonality of semantic and forgery subspaces, where the forgery pattern learning is conducted within a new subspace that is orthogonal to the original semantic subspace, avoiding the distortion of the pre-existing semantic knowledge. To achieve this, we introduce a novel approach that uses SVD to explicitly construct the orthogonal semantic and forgery subspaces by freezing the principal components for representing semantic knowledge, while adapting only the residual components for forgery pattern learning.
3.2 Effort Approach
The overall pipeline of the proposed Effort approach is illustrated in Fig. 4, aiming to address the asymmetry phenomena in AIGI detection. In general, our approach involves the Singular Value Decomposition (SVD) to construct explicit orthogonal semantic and forgery subspaces, avoiding the distortion of well-learned semantic knowledge during fake pattern learning.
Formally, given a pre-trained weight matrix $W\in\mathbb{R}^{d_{1}\times d_{2}}$ for a certain linear layer, we perform SVD to decompose $W$:
$W=U\Sigma V^{\top},$ (1)
where $U\in\mathbb{R}^{d_{1}\times d_{1}}$ and $V\in\mathbb{R}^{d_{2}\times d_{2}}$ are orthogonal matrices containing the left and right singular vectors, respectively, and $\Sigma\in\mathbb{R}^{d_{1}\times d_{2}}$ is a diagonal matrix with singular values on the diagonal. Since the linear layer of VFM generally has the same input and output dimensions, we consider the case of SVD with $d_{1}=d_{2}=n$ in the following discussion.
To obtain a rank-$r$ approximation of the pre-trained weight matrix, we retain only the top $r$ singular values and corresponding singular vectors:
$W\approx W_{r}=U_{r}\Sigma_{r}V_{r}^{\top},$ (2)
where $U_{r}\in\mathbb{R}^{n\times r}$, $\Sigma_{r}\in\mathbb{R}^{r\times r}$, and $V_{r}\in\mathbb{R}^{n\times r}$. We keep $W_{r}$ frozen during training to preserve the knowledge learned from real images.
The residual component, defined as the difference between the pre-trained weights and the SVD approximation, is used to learn representations specific to fake images:
$\Delta W=W-W_{r}=U_{n-r}\Sigma_{n-r}V_{n-r}^{\top},$ (3)
where $U_{n-r}\in\mathbb{R}^{n\times(n-r)}$, $\Sigma_{n-r}\in\mathbb{R}^{(n-r)\times(n-r)}$, $V_{n-r}\in\mathbb{R}^{n\times(n-r)}$. It is important to note that $\Delta W$ represents a learnable form associated with the remaining singular value decomposition, reflecting slight modifications or perturbations to the original weight matrix.
During training, we only optimize $\Delta W$ while keeping $U_{r}$, $\Sigma_{r}$, and $V_{r}$ fixed. This implementation ensures that the model retains its capability to process real images via the SVD approximation and adapts to detect deepfakes through the trivial residual components of the weight matrix.
To encourage the $\Delta W$ to capture both useful and meaningful discrepancy between the real and fake, it’s significant to guarantee that optimizing $\Delta W$ does not change the properties of the overall weight matrix $W$ (i.e., Minimize the impact on the real information of the pre-trained weight as much as possible). Thus, we proposed two constraints to realize this goal:
Orthogonal Constraint. We maintain the orthogonality among each singular vector to keep orthogonal subspace for learning real/fake:
$\mathcal{L}_{\rm orth}=\lVert\hat{U}^{\top}\hat{U}-I\rVert_{F}^{2}+\lVert\hat{V}^{\top}\hat{V}-I\rVert_{F}^{2},$ (4)
where $\hat{U}\in\mathbb{R}^{n\times n}$ denote the concatenation of $U_{r}$ and $U_{n-r}$ along the row dimension, $\hat{V}\in\mathbb{R}^{n\times n}$ denote the concatenation of $V_{r}$ and $V_{n-r}$ along the row dimension, $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm, and $I$ is the identity matrix of appropriate dimensions.
Singular Value Constraint. The singular values can be interpreted as a type of scaling that affects the magnitude of the corresponding singular vectors. There is a relationship between singular values and the Frobenius norm of the weight matrix being decomposed:
$\lVert W\rVert_{F}=\sqrt{\sum_{i}\sigma_{i}^{2}},$ (5)
where $\sigma_{i}$ denotes the $i$-th singular value of the corresponding weight matrix.
To maximize the reduction of the impact of real knowledge, we constrain the singular values of the optimized weight matrix $\hat{W}$ to remain consistent with those of the original weight matrix $W$:
$\mathcal{L}_{\rm ksv}=\bigg{|}\sum_{i=r+1}^{n}\hat{\sigma}_{i}^{2}-\sum_{i=r+1}^{n}\sigma_{i}^{2}\bigg{|}=\bigg{|}\lVert\hat{W}\rVert_{F}^{2}-\lVert W\rVert_{F}^{2}\bigg{|}$ (6)
[…]
Loss Function. The overall loss function for training the model combines the classification loss $\mathcal{L}_{\rm cls}$ (e.g., cross-entropy loss for binary classification) and the orthogonality regularization loss:
$\mathcal{L}=\mathcal{L}_{\rm cls}+\lambda_{1}\frac{1}{m}\sum_{i}^{m}\mathcal{L}_{\rm orth}^{i}+\lambda_{2}\frac{1}{m}\sum_{i}^{m}\mathcal{L}_{\rm ksv}^{i},$ (7)
where $\lambda_{1},\lambda_{2}$ are hyperparameters that balance the importance of the corresponding regularization term, and $m$ represents the number of pre-trained weight matrices on which our approach is applied.
In practice, we adapt our approach to the linear layers within the self-attention module across all transformer layers of the Vision Foundation Model (VFM) to leverage their rich, well-learned real distributions.
4 Experiment
Methods Trainable Cross-dataset Evaluation Cross-method Evaluation
Param. CDF-v2 DFD DFDC DFDCP DFo WDF FFIW Avg. UniFace BleFace MobSwap e4s FaceDan FSGAN InSwap SimSwap Avg.
F3Net [64] 22M 0.789 0.844 0.718 0.749 0.730 0.728 0.649 0.743 0.809 0.808 0.867 0.494 0.717 0.845 0.757 0.674 0.746
SPSL [47] 21M 0.799 0.871 0.724 0.770 0.723 0.702 0.794 0.769 0.747 0.748 0.885 0.514 0.666 0.812 0.643 0.665 0.710
SRM [53] 55M 0.840 0.885 0.695 0.728 0.722 0.702 0.794 0.767 0.749 0.704 0.779 0.704 0.659 0.772 0.793 0.694 0.732
CORE [58] 22M 0.809 0.882 0.721 0.720 0.765 0.724 0.710 0.762 0.871 0.843 0.959 0.679 0.774 0.958 0.855 0.724 0.833
RECCE [4] 48M 0.823 0.891 0.696 0.734 0.784 0.756 0.711 0.779 0.898 0.832 0.925 0.683 0.848 0.949 0.848 0.768 0.844
SLADD [8] 21M 0.837 0.904 0.772 0.756 0.800 0.690 0.683 0.777 0.878 0.882 0.954 0.765 0.825 0.943 0.879 0.794 0.865
SBI [72] 18M 0.886 0.827 0.717 0.848 0.899 0.703 0.866 0.821 0.724 0.891 0.952 0.750 0.594 0.803 0.712 0.701 0.766
UCF [91] 47M 0.837 0.867 0.742 0.770 0.808 0.774 0.697 0.785 0.831 0.827 0.950 0.731 0.862 0.937 0.809 0.647 0.824
IID [30] 66M 0.838 0.939 0.700 0.689 0.808 0.666 0.762 0.789 0.839 0.789 0.888 0.766 0.844 0.927 0.789 0.644 0.811
TALL [90] 87M 0.831 0.833 0.693 0.739 0.793 0.673 0.679 0.749 0.714 0.699 0.805 0.651 0.768 0.863 0.762 0.616 0.735
[…]
Effort (Ours) 0.19M 0.956 0.965 0.843 0.909 0.977 0.848 0.921 0.917 0.962 0.873 0.953 0.983 0.926 0.957 0.936 0.926 0.940
Table 1: Benchmarking Results of Cross-dataset Evaluations (Protocol-1) and Cross-method Evaluations (Protocol-2). All detectors are trained on FF++_c23 [70] and evaluated on other fake data. The best results are highlighted in bold and the second is underlined. $\dagger$ indicates the results are obtained by using the model’s checkpoint provided by the authors, otherwise, the results are cited from [92, 94, 11].
Methods GAN Deep fakes Low level Perceptual loss Guided LDM Glide Dalle mAP
Pro- GAN Cycle- GAN Big- GAN Style- GAN Gau- GAN Star- GAN SITD SAN CRN IMLE 200 steps 200 w/cfg 100 steps 100 27 50 27 100 10
CNN-Spot [86] 100.0 93.47 84.50 99.54 89.49 98.15 89.02 73.75 59.47 98.24 98.40 73.72 70.62 71.00 70.54 80.65 84.91 82.07 70.59 83.58
Patchfor [5] 80.88 72.84 71.66 85.75 65.99 69.25 76.55 76.19 76.34 74.52 68.52 75.03 87.10 86.72 86.40 85.37 83.73 78.38 75.67 77.73
Co-occurence [57] 99.74 80.95 50.61 98.63 53.11 67.99 59.14 68.98 60.42 73.06 87.21 70.20 91.21 89.02 92.39 89.32 88.35 82.79 80.96 78.11
Freq-spec [98] 55.39 100.0 75.08 55.11 66.08 100.0 45.18 47.46 57.12 53.61 50.98 57.72 77.72 77.25 76.47 68.58 64.58 61.92 67.77 66.21
[…]
UniFD [61] 100.0 98.13 94.46 86.66 99.25 99.53 91.67 78.54 67.54 83.12 91.06 79.24 95.81 79.77 95.93 93.93 95.12 94.59 88.45 90.14
[…]
NPR [79] 100.0 99.53 94.53 99.94 88.82 100.0 84.41 97.95 99.99 50.16 50.16 98.26 99.92 99.91 99.92 99.87 99.89 99.92 99.26 92.76
[…]
Effort (Ours) 100.0 100.0 99.99 99.77 100.0 100.0 98.95 97.53 97.53 100.0 100.0 95.39 99.99 99.89 100.0 99.87 99.92 99.98 99.96 99.41
Table 2: Benchmarking Results of Cross-method Evaluations in terms of AP Performance on the UniversalFakeDetect Dataset. $\dagger$ indicates that the results are obtained by using the official pre-trained model or reproduction.
Methods GAN Deep fakes Low level Perceptual loss Guided LDM Glide Dalle mAcc
Pro- GAN Cycle- GAN Big- GAN Style- GAN Gau- GAN Star- GAN SITD SAN CRN IMLE 200 steps 200 w/cfg 100 steps 100 27 50 27 100 10
CNN-Spot [86] 99.99 85.20 70.20 85.70 78.95 91.70 53.47 66.67 48.69 86.31 86.26 60.07 54.03 54.96 54.14 60.78 63.80 65.66 55.58 69.58
Patchfor [5] 75.03 68.97 68.47 79.16 64.23 63.94 75.54 75.14 75.28 72.33 55.30 67.41 76.50 76.10 75.77 74.81 73.28 68.52 67.91 71.24
Co-occurence [57] 97.70 63.15 53.75 92.50 51.10 54.70 57.10 63.06 55.85 65.65 65.80 60.50 70.70 70.55 71.00 70.25 69.60 69.90 67.55 66.86
Freq-spec 49.90 99.90 50.50 49.90 50.30 99.70 50.10 50.00 48.00 50.60 50.10 50.90 50.40 50.40 50.30 51.70 51.40 50.40 50.00 55.45
[…]
UniFD [61] 100.0 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.50 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
[…]
NPR [79] 99.84 95.00 87.55 96.23 86.57 99.75 76.89 66.94 98.63 50.00 50.00 84.55 97.65 98.00 98.20 96.25 97.15 97.35 87.15 87.56
[…]
Effort (Ours) 100.0 99.85 99.60 95.05 99.60 100.0 87.60 92.50 81.50 98.90 98.90 69.15 99.30 96.80 99.45 97.45 97.80 97.15 98.05 95.19
Table 3: Benchmarking Results of Cross-method Evaluations in terms of Acc Performance on the UniversalFakeDetect Dataset. $\dagger$ indicates that the results are obtained by using the official pre-trained model or reproduction.
4.1 Deepfake Image Detection
Implementation Details. We utilize CLIP ViT-L/14 [65] as the default vision foundation model (VFM). We also investigate other VFMs in Tab. 5. We follow the pre-processing and training pipeline and use the codebases of DeepfakeBench [92]. Additionally, we sample 8 frames from each video for training and 32 frames for inference, following [72]. We use the fixed learning rate of 2e-4 for training our approach and employ the Adam [36] for optimization. We set the batch size to 32 for both training and testing. We also employ several widely used data augmentations, such as Gaussian Blur and Image Compression, following other existing works [93, 72, 11]. For the evaluation metric, we report the widely-used video-level Area Under the Curve (AUC) to compare our approach with other works, following [46, 90, 72]. Similar to these works, we compute the average model’s output probabilities of each video to obtain the video-level AUC.
Evaluation Protocols and Dataset.
We adopt two widely used and standard protocols for evaluation: Protocol-1: cross-dataset evaluation and Protocol-2: cross-manipulation evaluation within the FF++ data domain. For Protocol-1, we conduct evaluations by training the models on FaceForensics++ (FF++) [70] and testing them on other seven deepfake detection datasets: Celeb-DF-v2 (CDF-v2) [45], DeepfakeDetection (DFD) [16], Deepfake Detection Challenge (DFDC) [15], the preview version of DFDC (DFDCP) [19], DeeperForensics (DFo) [32], WildDeepfake (WDF) [105], and FFIW [101]. Note that FF++ has three different compression versions and we adopt the c23 version for training all methods in our experiments, following most existing works [93]. For Protocol-2, we evaluate the models on the latest deepfake dataset DF40 [95], which contains the forgery data generated within the FF++ domain, ensuring the evaluation fake methods different but the data domain unchanged. In this manner, we can assess the model’s generalization when facing unseen fake methods, ignoring the impact of the data domain gap.
Evaluation Benchmarking.
To provide a comprehensive benchmark for comparison, we introduce 13 competitive detectors, including several classical detection methods such as F3Net [64] (ECCV’20), SPSL [47] (CVPR’20), SRM [47] (CVPR’21), CORE [58] (CVPRW’22), RECCE [4] (CVPR’22), and SBI [72] (CVPR’22), and also several latest SOTA methods (after 2023), such as UCF [91] (ICCV’23), IID [30] (CVPR’23), TALL [90] (ICCV’23), LSDA [94] (CVPR’24), ProDet [11] (NeurIPS’24), and CDFA [46] (ECCV’24). All detectors are trained on FF++ (c23) and tested on other fake data. Results in Tab. 1 demonstrate two notable advantages of our approach. (1) generalizability: we see that our approach consistently and largely outperforms other models across basically all tested scenarios, validating that our method is generalizable for detecting unseen fake data, even for the latest face-swapping techniques such as BleFace [73]. (2) efficiency: it is worth noting that our method only needs 0.19M parameters for training to achieve superior generalization. As we can see most latest SOTA detectors such as LSDA and ProDet all use about 100M parameters for training, while we are about 1,000$\times$ smaller.
4.2 Synthetic Image Detection
Evaluation Metrics.
We follow existing works [85, 61, 48] for benchmarking and report both average precision (AP) and classification accuracy (Acc). For Acc, we set the classification threshold for each dataset to $0.5$ to ensure a fair comparison.
Table 4: Ablation studies regarding the tunable $n-r$ values in SVD. All models are trained on FF++ (c23) and tested on CDF-v2 and SimSwap. “FFT” indicates the full fine-tuning. “Linear-Prob” indicates fine-tuning the FC layer only.
Archs. #Params $n-r$ CDF-v2 SimSwap Avg.
Baseline (Linear-Prob) 0.002M 0.765 0.769 0.767
Baseline (FFT) 307M 0.857 0.860 0.859
Ours Variant-1 12M 64 0.946 0.935 0.941
Variant-2 3M 16 0.948 0.930 0.939
Variant-3 0.75M 4 0.955 0.921 0.938
Variant-4 0.19M 1 0.956 0.926 0.941
Table 5: Ablation studies regarding different vision foundation models (VFMs) were used. All models are trained on FF++ (c23) and tested on CDF-v2 and SimSwap.
VFMs #Params #ImgSize CDF-v2 SimSwap Avg.
BEIT-v2 [63] 303M 224 0.855 0.821 0.838
+ Ours 0.14M 224 0.894 0.850 0.872
SigLIP [96] 316M 256 0.877 0.713 0.795
+ Ours 0.19M 256 0.895 0.778 0.867
CLIP [65] 307M 224 0.857 0.860 0.859
+ Ours 0.19M 224 0.956 0.926 0.941
UniversalFakeDetect Dateset.
We adhere to the protocol outlined in [85, 61] and utilize ProGAN’s real and fake images as our training dataset, which includes 20 subsets of generated images. The evaluation set contains 19 subsets derived from different kinds of generative models, including ProGAN [33], CycleGAN [102], BigGAN [2], StyleGAN [34], GauGAN [62], StarGAN [13], DeepFakes [69], SITD [7], SAN [14], CRN [10], IMLE [43], Guided (guided diffusion model)[17], LDM (latent diffusion model) [67], Glide [60], and DALLE [66].
Implementation Details.
Similar to the setting of deepfake image detection, we adopt pre-trained CLIP ViT-L/14 as the backbone and use the Adam optimizer [36] with a fixed learning rate of 2e-4. The batch size is set to 48. Other settings and details are the same with [61].
Evaluation Analysis.
The AP and Acc results are presented in Tab. 2 and Tab. 3, respectively. Our method attains impressive detection results, achieving 95.19% mAcc and 99.41% mAP across the 19 test subsets. One similar approach to ours is UniFD, which also preserves the original semantic knowledge of CLIP and fine-tunes only the FC layer for discrimination. In contrast to UniFD which performs discrimination in the semantic space, our approach utilizes SVD to create an orthogonal low-ranked subspace for learning forgeries while preserving the essential high-ranked semantic space, achieving the discrimination by leveraging both semantic and forgery subspaces, achieving a significant improvement of 9.27% in mAcc and 13.81% in mAP over UniFD. Besides, when compared to the SoTA method, FatFormer, we achieve 4.33% mAcc improvement without relying on the extra text encoder of CLIP. This further demonstrates the superiority of our approach.
Table 6: Ablation studies regarding the singular value constraint ($\mathcal{L}_{\rm ksv}$) and orthogonal constraint ($\mathcal{L}_{\rm orth}$). All models are trained on FF++ (c23) and tested on other datasets.
Ours CDF-v2 DFDC SimSwap FSGAN Avg.
SVD $\mathcal{L}_{\rm ksv}$ $\mathcal{L}_{\rm orth}$
$\times$ $\times$ $\times$ 0.857 0.758 0.860 0.939 0.854
$\times$ $\times$ 0.940 0.829 0.910 0.955 0.909
$\times$ 0.944 0.841 0.927 0.953 0.916
$\times$ 0.945 0.847 0.914 0.950 0.914
0.956 0.843 0.926 0.957 0.921
4.3 Ablation Study and Analysis
How many singular values should be tuned when applying SVD?
When applying SVD, we create the semantic (frozen) and forgery (tuning) subspaces. But a crucial question arises: how many singular values should be set aside for freezing and how many for tuning? To figure out the optimal empirical values for this, we carry out an ablation study. The purpose of this study is to evaluate the effectiveness of different values of $n-r$. Here, $n$ represents the full-ranked dimension of the original weights, while $r$ is the low-ranked dimension that we intend to tune We examine four different variants with specific values for $n-r$, namely 64, 16, 6, and 1. Additionally, we include the baseline case, which is the full-finetune CLIP, for the sake of comparison. Results in Tab. 4 suggest that our approach largely outperforms the baseline. Moreover, the performance levels of the four variants are quite similar. This implies that none of these variants holds a distinct advantage when it comes to generalization. However, among them, variant-4 has the fewest trainable parameters. Consequently, we select this particular setting as the default option in our other experiments.
Figure 5: Evidence for verifying the illustration of Fig. 2, showing the latent distributions of vanilla CNN (e.g., Res50 [86]) and ours.
Figure 6: Evidence for validating the discrimination dimension from the logits space. Please refer to the text for details.
Compatibility with other vision foundation models.
By default, we choose CLIP as the vision foundation model (VFM) in our experiments. To validate the generality and versatility of our approach, we conduct an ablation study to apply our method to other VFMs, including BEIT-v2 [63] and SigLIP [96]. Results in Tab. 5 show that our approach can be seamlessly applied to other existing VFMs for improving the model’s generalization performance.
Figure 7: Evidence for validating the discrimination dimension from the feature space. We use PCA to analyze the rank of the feature space. Please refer to the text for details.
How to verify the naively trained models use only forgery patterns for discrimination?
In our motivation (RQ-2), we argue that the naively trained models rely solely on forgery patterns for discrimination, as the models tend to overfit the seen fake patterns during training (see Fig. 1, Fig. 3, Fig. 5). To verify that the baseline models (a vanilla CNN, i.e., Xception) indeed leverages only the forgery patterns for discrimination, we conduct a new visualization in Fig. 6, providing reasonable evidence to see the model’s discrimination behavior. Specifically, we visualize the model’s output, i.e., logits, for both real and fake classes, and a dotted line for the boundary (to those samples with real logits equal to fake logits). As shown in Fig. 6, we see that all prediction samples of Xception lie in a linear line (similar to $y=-x+b$), indicating that the Xception relies on one dimension for discriminating real and fake. In addition, the CLIP contains rich semantic knowledge, but fully fine-tuning CLIP can distort the well-learned semantic knowledge, making it near to a line ($y=-x+b$). Finally, our approach maximizes the retention of semantic knowledge. We achieve this by learning forgeries within a new orthogonal subspace relative to the semantic subspaces, achieving superior generalization than the other two models.
Low-ranked feature space causes the model’s overfitting.
We use Principal Component Analysis (PCA) to visualize the feature space with the two most principal components (the top row) and demonstrate the explained variance ratio of different principal components (the below row), which is similar to “how much of the total information of the data is captured by each principal component”. Results in Fig. 7 show that the feature space of the Xception is highly low-ranked, with only two principal components to capture all information, resulting in limited generalization.
5 Conclusion
In this paper, we provide a new perspective to explain the failure reason of the generalization in AIGI detection, namely the asymmetry phenomena, where a naively trained detector very quickly shortcuts to the seen fake patterns, dominating the whole feature space and shaping it to be highly low-ranked, thereby limiting its expressivity and generalization. The key idea of this paper is to incorporate the rich semantic knowledge within the vision foundation models (VFMs) to expand the discrimination and feature space to higher-ranked, alleviating the model’s overfitting to seen fake patterns. A straightforward strategy is to directly fine-tune the VFM, but it risks distorting the learned semantic space, pushing the model back toward overfitting. To this end, we propose to decompose the original semantic space into two orthogonal subspaces for preserving semantic knowledge while learning forgery. We conduct extensive experiments on both deepfake and synthetic image detection benchmarks, showing the superior advantages in both generalization and efficiency.
Appendix
This appendix material provides additional experimental results, visualizations, and future works:
  • Algorithm procedure of the proposed approach (see Sec. A).
  • Additional experimental results and ablations (see Sec. B).
  • Additional visualizations and analysis (see Sec. D).
  • Future works and limitation discussion (see Sec. E).
Appendix A Algorithm Procedure Illustration
Here, we provide an algorithm illustration of the proposed Effort approach in Alg. 1 for a comprehensive understanding.
Algorithm 1 Effort Approach Algorithm
0:  Pre-trained weight matrix $W\in\mathbb{R}^{n\times n}$; Rank $r$; Training data $\mathcal{D}=\{(x_{i},y_{i})\}^{N}_{i=1}$; Hyperparameters $\lambda_{1}$, $\lambda_{2}$
0:  Updated weight matrix $W$
1:  $\triangleright$ Step 1: Singular Value Decomposition
2:  Decompose $W$ via SVD: $W=U\Sigma V^{\top}$
3:  Retain top $r$ singular values and vectors:
4:   $U_{r}\in\mathbb{R}^{n\times r}$, $\Sigma_{r}\in\mathbb{R}^{r\times r}$, $V_{r}\in\mathbb{R}^{n\times r}$
5:  Compute $W_{r}=U_{r}\Sigma_{r}V_{r}^{\top}$
6:  Keep $W_{r}$ fixed during training
7:  Compute residual component: $\Delta W=W-W_{r}$
8:  Decompose $\Delta W$ via SVD: $\Delta W=U_{n-r}\Sigma_{n-r}V_{n-r}^{\top}$
9:  Initialize $\Delta W$
10:  Define concatenated matrices:
11:   $\hat{U}=[U_{r},U_{n-r}]\in\mathbb{R}^{n\times n}$
12:   $\hat{V}=[V_{r},V_{n-r}]\in\mathbb{R}^{n\times n}$
13:  $\triangleright$ Step 2: Training Loop
14:  for each epoch do
15:     for each batch in $\mathcal{D}$ do
16:        $\triangleright$ Forward Pass
17:        Compute model output using $W=W_{r}+\Delta W$
18:        Compute classification loss $\mathcal{L}_{\rm cls}$
19:        $\triangleright$ Compute Constraints
20:        Compute orthogonality loss:
21:         $\mathcal{L}_{\rm orth}=\left\|\hat{U}^{\top}\hat{U}-I\right\|_{F}^{2}+\left\|\hat{V}^{\top}\hat{V}-I\right\|_{F}^{2}$
22:        Compute singular value constraint loss:
23:         $\mathcal{L}_{\rm ksv}=\left|\left\|\hat{W}\right\|_{F}^{2}-\left\|W\right\|_{F}^{2}\right|$
24:        $\triangleright$ Total Loss
25:        $\mathcal{L}=\mathcal{L}_{\rm cls}+\lambda_{1}\mathcal{L}_{\rm orth}+\lambda_{2}\mathcal{L}_{\rm ksv}$
26:        $\triangleright$ Backward Pass and Optimization
27:        Update $\Delta W$ using gradient descent to minimize $\mathcal{L}$
28:     end for
29:  end for
30:  $\triangleright$ Step 3: Output
31:  Update the weight matrix: $W\leftarrow W_{r}+\Delta W$
32:  return  Updated weight matrix $W$
Appendix B Additional Results and Ablations
In this section, we provide additional experimental results and ablation studies to our proposed approach.
B.1 Results on GenImage Benchmark
In our manuscript, we present the benchmarking outcomes of the UniversalFakeDetect Dataset. Additionally, we report the results obtained from another widely utilized benchmark known as GenImage [103]. This GenImage dataset predominantly utilizes the Diffusion model for image generation, incorporating models such as Midjourney [56], SDv1.4 [68], SDv1.5 [68], ADM [18], GLIDE [59], Wukong [1], VQDM [22], and BigGAN [3]. Following the settings defined for GenImage, we designate SDv1.4 as the training set and the remaining models as the test set. Given the diverse image sizes within the GenImage dataset, images with a size smaller than 224 pixels are duplicated and subsequently cropped to 224 pixels, following [79]. We employ the same setting to re-implement FreqNet, FatFormer, and NPR, and also report the results of UnivFD and DRCT from [6].
The results on the GenImage dataset are presented in Table 7. When SDv1.4 is employed as the training set, our method attains an overall accuracy rate of 91.1% across the entire test set. Compared to similar methods that utilize CLIP as the backbone, such as UnivFD and FatFormer, our approach improves accuracy by 11.6% and 2.2%, respectively. Moreover, when contrasted with the latest state-of-the-art (SOTA) method DRCT (ICML 2024), the proposed method achieves a 1.6% enhancement in accuracy. This clearly indicates that our method demonstrates superior generalization capabilities and achieves SOTA performance on the GenImage benchmark.
Methods Venues Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN mAcc
ResNet-50 [27] CVPR 2016 54.9 99.9 99.7 53.5 61.9 98.2 56.6 52.0 72.1
DeiT-S [82] ICML 2021 55.6 99.9 99.8 49.8 58.1 98.9 56.9 53.5 71.6
Swin-T [49] ICCV 2021 62.1 99.9 99.8 49.8 67.6 99.1 62.3 57.6 74.8
CNNSpot [86] CVPR 2020 52.8 96.3 95.9 50.1 39.8 78.6 53.4 46.8 64.2
Spec [98] WIFS 2019 52.0 99.4 99.2 49.7 49.8 94.8 55.6 49.8 68.8
F3Net [64] ECCV 2020 50.1 99.9 99.9 49.9 50.0 99.9 49.9 49.9 68.7
GramNet [50] CVPR 2020 54.2 99.2 99.1 50.3 54.6 98.9 50.8 51.7 69.9
UnivFD [61] CVPR 2023 91.5 96.4 96.1 58.1 73.4 94.5 67.8 57.7 79.5
NPR  [79] CVPR 2024 81.0 98.2 97.9 76.9 89.8 96.9 84.1 84.2 88.6
FreqNet  [78] AAAI 2024 89.6 98.8 98.6 66.8 86.5 97.3 75.8 81.4 86.8
FatFormer [48] CVPR 2024 92.7 100.0 99.9 75.9 88.0 99.9 98.8 55.8 88.9
DRCT [6] ICML 2024 91.5 95.0 94.4 79.4 89.2 94.7 90.0 81.7 89.5
Ours 82.4 99.8 99.8 78.7 93.3 97.4 91.7 77.6 91.1
Table 7: Benchmarking results of cross-method evaluations in terms of Acc performance on the Genimage dataset. We follow [103] and use the SDv1.4 as the training set while others as the testing sets. We directly cite the results of ResNet-50, DeiT-S, Swin-T, CNNSpot, Spec, F3Net, and GramNet from [103]. We obtain the results of UnivFD and DRCT from [6], and FreqNet, NPR, and FatFormer by using the official checkpoints for reproduction. We report the Accuracy metric for comparison following [6].
Methods Venues CDF-v2 DFDC Avg.
LipForensics [25] CVPR 2021 0.824 0.735 0.780
FTCN [100] ICCV 2021 0.869 0.740 0.805
HCIL [23] ECCV 2022 0.790 0.692 0.741
RealForensics [26] CVPR 2022 0.857 0.759 0.808
LTTD [24] NeurIPS 2022 0.893 0.804 0.849
AltFreezing [88] CVPR 2023 0.895
TALL-Swin [90] ICCV 2023 0.908 0.768 0.838
StyleDFD [12] CVPR 2024 0.890
NACO [97] ECCV 2024 0.895 0.767 0.831
Ours 0.956 0.843 0.900
Table 8: Cross-dataset generalization evaluations with existing SOTA video detectors. The results of other detectors are directly cited from their original papers. The metric is video-level AUC.
B.2 Comparison with Existing Video Detectors
In the manuscript, we mainly compare our method with image detectors. Here, we provide an individual result to compare our approach with existing SOTA video detectors. Following [97, 90], we conduct evaluations on the widely-used CDF-v2 [45] and DFDC [15] using the video-level AUC metric. We have considered both the classical detectors such as LipForensics and the latest SOTA detectors such as NACO (ECCV’24) for a comprehensive comparison. Results in Tab. 8 demonstrate that our image-based approach achieves higher generalization performance in both CDF-v2 and DFDC, improving 4.5% and 3.9% points than the second-best video-based models. This further validates the superior generalization performance of our approach.
Methods CDF-v2 DFD DFDC Avg.
LoRA [37] 0.838 0.834 0.717 0.796
MoE-LoRA [38] 0.867 0.904
Dual-Adapter [71] 0.717 0.727
Ours 0.901 0.923 0.798 0.874
Table 9: Cross-dataset generalization evaluations with other adapter-based partially fine-tuned detectors. The results are cited from their original papers. The metric is frame-level AUC.
Archs. $n-r$ $r$ mAcc
UniFD [61] (Linear-Prob) 81.02
Baseline (FFT) 86.22
Ours Variant-1 256 92.13
Variant-2 64 93.68
Variant-3 16 94.45
Variant-4 4 94.37
Variant-5 1 95.19
LoRA Variant-1 256 91.42
Variant-2 64 91.06
Variant-3 16 91.89
Variant-4 4 93.53
Variant-5 1 93.03
Table 10: Ablation studies on synthetic image detection (UniversalFakeDetect Dataset) regarding the tunable $n-r$ values in SVD (Ours) and $r$ values in LoRA. All models are trained on ProGAN’s images and tested on 19 different kinds of generative models’ images. “FFT” indicates the full fine-tuning. “Linear-Prob” indicates fine-tuning FC layer only, where we reproduce the results from UniFD [61].
B.3 Within-Domain Evaluation Results
In the previous contents of the manuscript, we mainly show the cross-domain generalization results. For a more comprehensive evaluation, we provide our method’s within-domain evaluations. Since we use FF++ [70] (c23) for training, following DeepfakeBench [92], we train our model on FF++ (c23) and test it on FF++ (c23) and FF++ (c40), where c40 is a compression version of c23. We compare the results of our approach with other competing baseline models such as CNN-Aug, F3Net, SPSL, SRM, and RECCE from DeepfakeBench. We also compare our approach with the latest SOTA detector, i.e., CDFA [46] (ECCV’24) for a comprehensive comparison. Results in Tab. 11 show that generally, all detectors perform very well on FF++ (c23), as the training and testing distributions are similar. However, the performance largely drops when applying compression (tested on the c40 version). Also, our approach demonstrates the best results on FF++ (c40), showing the robustness of our detector. Notably, it can be seen that although CDFA achieves superior generalization results on cross-domain evaluation scenarios, it’s performance on within-domain evaluation is limited. In contrast, our approach shows high results on both within-domain and cross-domain evaluations. Regarding the model’s robustness toward the post-processing, we provide another robustness experiment following [32] (refer to Sec. B.5 for details).
B.4 Comparison with Existing Adapter-Based Detectors
Parameter-efficient fine-tuning (PEFT) has become a popular technique for adapting pre-trained vision foundation models (VFMs) to downstream tasks. Low-ranked adaptation (LoRA) [29] is a widely-used approach for PEFT. Previous works [37, 48] employing LoRA in the VFMs have achieved good empirical generalization results on both deepfake detection and synthetic image detection benchmarks. Compared to full parameter fine-tuning (FFT), LoRA also has the advantage of retaining part of the original well-learned pre-trained semantic knowledge while learning the forgery patterns. Additionally, [38] and [71] introduce MoE-LoRA techniques and dual adapters into the deepfake detection fields. However, our SVD-based approach differs from these adapter-based methods and shows several advantages over them. Specifically, our method explicitly constructs two orthognoal subspaces for semantic and forgery using SVD, ensuring the pre-existing semantic knowledge will not be distorted and well-retained during the process of learning forgery. In contrast, the mentioned adapter-based methods do not explicitly ensure the orthogonality between semantic and forgery, still having the potential to distort the pre-existing semantic knowledge and result in unexpected generalization results.
To show our approach can achieve better generalization results than other adapter-based methods in both deepfake detection and synthetic image detection benchmarks, we provide several empirical results in Tab. 9 and Tab. 10. From these results, we can see that our proposed SVD-based method achieves clearly higher generalization results than adapter-based methods in both deepfake detection and synthetic image detection fields, as our approach explicitly preserves the semantic knowledge while learning the forgery patterns.
Methods FF-c23 FF-c40 Avg.
CNN-Aug [86] 0.8493 0.7846 0.8170
F3Net [64] 0.9635 0.8271 0.8953
SPSL [47] 0.9610 0.8174 0.8892
SRM [53] 0.9576 0.8114 0.8845
RECCE [4] 0.9621 0.8190 0.8906
[…]
Ours 0.9872 0.8429 0.9151
Table 11: Within-domain evaluations using the video-level AUC metric. All detectors are trained on FF-c23 and evaluated on other data. $\dagger$ donates our reproduction using the pre-trained weights, otherwise, we obtain the results from [92].
Methods GID-DF GID-F2F
Acc AUC Acc AUC
EfficientNet [80] 82.40 91.11 63.32 80.10
MLGD [41] 84.21 91.82 63.46 77.10
LTW [74] 85.60 92.70 65.60 80.20
DCL [75] 87.70 94.90 68.40 82.93
M2TR [84] 81.07 94.91 55.71 76.99
F3Net [64] 83.57 94.95 61.07 81.20
F2Trans [55] 89.64 97.47 81.43 90.55
CFM [52] 85.00 92.74 76.07 84.55
FA-ViT [51] 92.86 98.10 82.57 91.20
Ours 95.71 99.26 85.71 93.83
Table 12: Additional results of cross-manipulation evaluation on FF++ (c23). Following [55, 51], we conduct evaluations by training on the other three manipulated methods while testing on the remaining one. Specifically, GID-DF means training on the other three manipulated methods (FF-F2F, FF-FS, FF-NT) while testing on the FF-DF. Results of other methods are cited from [55, 51].
B.5 Robustness Evaluation
To evaluate our model’s robustness to random perturbations, we adopt the methodology used in previous studies [25, 100], which involves examining three distinct types of degradation: Block-wise distortion, Change contrast, and JPEG compression. We apply each of these perturbations at five different levels to assess the model’s robustness under varying degrees of distortion, following [9, 93]. The video-level AUC results for these unseen perturbations, using the model trained on FF++ (c23), are depicted in Fig. 8. Generally, our approach shows higher results than other methods, demonstrating the better robustness of our approach than other models.
Figure 8: Robustness to unseen perturbations. We present video-level AUC for five distinct degradation levels across three types of perturbations in [32].
Appendix C Additional cross-method evaluation
The capability of face forgery detectors to generalize to new manipulation methods is crucial in practical, real-world applications. In our manuscript, we present cross-method evaluations using the DF40 dataset [95]. Specifically, we train the models with four manipulation methods from FF++ (c23) and then test them on the other eight manipulation techniques provided in DF40. Furthermore, we conduct an additional cross-method evaluation following the protocol introduced in [75, 55, 51]. This protocol involves training the model on diverse manipulation types of samples and subsequently testing it on unknown manipulation methods. The results of this evaluation are reported in Tab. 12. It is evident that our proposed method attains remarkable performance in cross-manipulation evaluation. In terms of accuracy (ACC), it outperforms the latest SOTA detector FA-ViT by 2.85% on GID-DF and 3.14% on GID-F2F, respectively.
Figure 9: Logits and confidence distributions of the vanilla CNN (Xception) and ours. We show that (a) the Xception fits the fake patterns in the training set well (the left column), while other information is not adequately captured, leading to poor generalization when detecting unseen deepfake datasets (the right column). In contrast, our approach (b) learns both forgery patterns and semantic knowledge well, achieving more robust generalization performance.
VFMs #Params #ImgSize CDF-v2 SimSwap Avg.
BEIT-v2 [63] 303M 224 0.855 0.821 0.838
+ Ours 0.14M 224 0.894 0.850 0.872
SigLIP [96] 316M 256 0.877 0.713 0.795
+ Ours 0.19M 256 0.895 0.778 0.867
CLIP [65] 307M 224 0.857 0.860 0.859
+ Ours 0.19M 224 0.956 0.926 0.941
Table 13: Ablation studies on deepfake image detection (Cross-dataset) regarding different vision foundation models (VFMs) were used. All models are trained on FF++ (c23) and tested on CDF-v2 and SimSwap.
VFMs #Params #ImgSize mAP mAcc
BEIT-v2 [63] 303M 224 93.50 79.11
+ Ours 0.14M 224 97.39 83.66
SigLIP [96] 316M 256 94.30 81.23
+ Ours 0.19M 256 96.24 90.46
CLIP [65] 307M 224 97.95 86.22
+ Ours 0.19M 224 99.41 95.19
Table 14: Ablation studies on synthetic image detection (UniversalFakeDetect Dataset) regarding different vision foundation models (VFMs) were used. All models are trained on ProGAN’s images and tested on 19 different generative models’ images.
C.1 Additional Ablation Studies
Impact of Different Vision Foundation Models
We initialize the ViT backbone with several widely used pre-trained weights from different vision foundation models, including BEIT-v2 [63], CLIP [65], and SigLIP [96]. The results are shown in Tab. 13 and Tab. 14. It is evident that our proposed approach improves the generalization performance of different pre-trained ViTs. On the other hand, we note that different initialization significantly impacts generalization performance, indicating the importance of choosing a suitable pre-trained initialization. Through empirical results, we discover that the ViT pre-trained on CLIP exhibits the highest performance in both deepfake detection and synthetic image detection tasks. Therefore, we choose CLIP as the default setting for our approach.
VFMs #Params #ImgSize CDF-v2 SimSwap Avg.
CLIP-Base/16 86M 224 0.854 0.833 0.844
+ Ours 0.07M 224 0.915 0.919 0.917
CLIP-Large/14 307M 224 0.857 0.860 0.859
+ Ours 0.19M 224 0.956 0.926 0.941
Table 15: Ablation studies on deepfake image detection (Cross-dataset) regarding different ViT architectures were used. We employ the two architectures implemented in the original paper of CLIP [65] for experiments. All models are trained on FF++ (c23).
VFMs #Params #ImgSize mAP mAcc
CLIP-Base/16 86M 224 96.25 82.52
+ Ours 0.07M 224 98.47 88.46
CLIP-Large/14 307M 224 97.95 86.22
+ Ours 0.19M 224 99.41 95.19
Table 16: Ablation studies on synthetic image detection (UniversalFakeDetect Dataset) regarding different architectures were used. All models are trained on ProGAN’s images and tested on 19 different kinds of generative models’ images.
Impact of Different ViT Backbones
Here, we investigate the effects of different ViT architectures. Specifically, we consider two backbones that were implemented in the original paper of CLIP: ViT-Base-16 and ViT-Large-14. We conduct evaluations on both deepfake detection and synthetic image detection benchmarks, as shown in Tab. 15 and Tab. 16. Compared to fully fine-tuning the CLIP model, our proposed approach consistently demonstrates substantial enhancements in generalization performance across these ViT backbones. It is worth noting that CLIP-Large performs better than CLIP-Base by a notable margin. Based on this ablation experiment, we ultimately choose ViT-Large as our default backbone.
Appendix D Additional Analysis and Visualizations
D.1 Logit Distribution Analysis
We show the prediction distribution of a naive trained detector’s output (real probabilities) in Fig. 9. We see that the vanilla CNN detector tends to learn the fake in the training set and cannot learn a good feature of real when evaluating unseen fakes. When the fake patterns are unseen during training, the logit for real/fake classification tends to be very unconfident. Additionally, the overlapping between real and fake classes of Xception is significantly increased when detecting unseen the forgery dataset. In contrast, our approach preserves semantic knowledge while learning forgery patterns during training, thereby improving the model’s generalization performance.
Figure 10: t-SNE visualizations of the latent feature spaces between vanilla CNN [86] and ours. We use the testing set of ProGAN and StyleGAN within UniversalFakeDetect Dataset [86] for visualization.
Figure 11: t-SNE visualizations of the latent feature spaces between vanilla CNN [86] and ours. We use the testing set of CycleGAN within UniversalFakeDetect Dataset [86] for visualization.
D.2 Additional t-SNE Visualizations
In our manuscript, we visualize the t-SNE of the seen fake ProGAN and unseen fake StyleGAN for the comparison of vanilla CNN (Res-50 [86]) and ours (see Fig. 10). Now, we provide additional t-SNE visualizations, including t-SNE results between vanilla CNN and ours on CycleGAN, as shown in Fig. 11. As we can see from both two figures, our approach maximizes and preserves the semantic knowledge while fitting the forgery patterns during training, whereas the vanilla CNN overfits the seen fake method, learning forgery patterns only, thereby resulting in a highly low-ranked feature space (see Fig. 6 and Fig. 7 of the manuscript for details) and causing the overfitting to seen forgery patterns in the training set. Additionally, we see that the logit distribution of the vanilla CNN has a larger overlapping between fake and real, while ours is highly smaller, suggesting that our approach achieves a better generalization performance.
Appendix E Limitation and Future Work
The core idea of this paper is to decompose the original semantic space into two orthogonal subspaces for preserving semantic knowledge while learning the forgery. In our manuscript and appendix, we have conducted extensive experiments and in-depth analysis on both deepfake and synthetic image detection benchmarks, showing the superior advantages in both generalization and efficiency. One limitation of our work is that our approach regards all fake methods in one class during training real/fake classifiers, potentially ignoring the specificity and generality of different fake methods.
In the future, we plan to expand our approach into a incremental learning framework, where each fake method will be regarded as “one SVD branch”, ensuring the orthogonality between different fake methods, thereby avoiding the severe forgetting of previous learned fake methods. This extension design will help our approach better address the future deepfake types in the real-world scenario. Additionally, although our work’s scope mainly focuses on deepfake and synthetic image detection, our approach also has the potential to be applied to other similar fields such as face anti-spoofing, anomaly detection, etc. Furthermore, we hope our proposed approach can inspire future research in developing better orthogonal modeling strategies.
Ethics & Reproducibility. All of the facial images that are utilized are sourced from publicly available datasets and are accompanied by appropriate citations. This guarantees that there is no infringement upon personal privacy. We will make all codes and checkpoints available for public access upon acceptance.